OpenStudy (anonymous):

for f(x)= 4x+1 and g(x)= x^2-5, find (f*g)(x)

OpenStudy (trojanpoem):

Normal multiplication ; f(x *g)(x) = [4x+1] * [x^2 -5] = 4x^3 - 20 x + x^2 -5 = 4x^3 + x^2 - 20x - 5

OpenStudy (trojanpoem):

Notice that f(x*g)(x) = f(x) * g(x) , f(x/g) = f(x)/ g(x) f(x -g ) = f(x) _ g(x) , f(x+g) = f(x) + g(x)

OpenStudy (anonymous):

so f(x)=2x+1 and g(x)=x^2-7, find (f-g)(x) it will be 2x^2-15

OpenStudy (trojanpoem):

(f-g)(x) =2x +1 -x^2 -7 = 2x - x^2 -6